OpenStudy (anonymous):

if cot(B)=3 and cot=cos/sin. then is cos=3 and sin=1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay i did something wrong though

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

tan=1/3 but i got sin, cos, sec, csc wrong

OpenStudy (anonymous):

i said sin=1, cos=3, sec=1/3 and csc=1

OpenStudy (anonymous):

sec B= 1/cosB = 1/3 csc B = 1/ sin B = 1/1 = 1

myininaya (myininaya):

if cot(B)=3/1 then we have the oppsite side of B is 1 and the adjacent side to B is 3 heres a triangle